In a triangle ABC, angle B is 4 times angle A and C is 19 less than 5 times angle A
Answer by mananth(16946)

Let A be x deg
angle B will be 4x
angle c= 5x-19
..
x+4x+5x-19=180
10x=199
x=19.9 deg which is angle C
Angle B = 4x=4*19.9=79.6 deg
Angle C = 5x-19= 5*19.9-19
99.5-19= 80.5 deg